# Corresponding author(s)

### v1

@@ 4.3 Reasoning Capability Evaluation
Figure 13: Comparison of DeepSeek-R1 distilled models and safety enhancement models on reasoning benchmarks.
We further evaluate the reasoning capabilities of the safety-enhanced models to investigate whether our safety enhancement methodology compromises their reasoning performance. Figure 13 presents a comparative analysis of reasoning capabilities before and after safety enhancement. As evidenced by the results, safety enhancement does not significantly impair the reasoning abilities of the models; on the contrary, it concurrently enhances safety while improving the reasoning performance of certain models. The distilled models exhibit varied performance on two mathematical reasoning benchmarks: In the MATH-500 test, DeepSeek-R1-Distill-Llama-8B shows a maximum decline of 3.1% in reasoning capability, while DeepSeek-R1-Distill-Qwen-14B achieves a maximum improvement of 1.1%; in the GPQA test, the reasoning capability of DeepSeek-R1-Distill-Llama-70B shows a maximum decrease of 1.6%, while DeepSeek-R1-Distill-Qwen-14B demonstrates a maximum increase of 5.0%. Furthermore, in the code benchmark test, reasoning capabilities exhibit a marginal reduction, with the decline ranging from 0.3% to 4.1%. Importantly, the observed reductions in code capability across all models are substantially outweighed by the gains in safety performance.
@@ 4.4 Overall Performance
Table 3 presents the overall performance of the base model series, the distilled model series, and the safety-enhanced model series on two Chinese safety evaluation tasks. The overall scores for each series are obtained by calculating the mean scores of the six models within the series. The results indicate that the distilled models exhibit a 2.17% decrease in accuracy on the risk content identification task compared to the non-distilled base models. On the refusal to answer task, the distilled models show reductions of 1.29%, 4.93%, and 2.45% in the RR-1, RR-2, and HR metrics, respectively. This suggests that while distillation enhances the models’ reasoning capabilities, it simultaneously weakens their ability to identify risky content and to reject risky queries while providing responsible responses. Although the harmful response rate is mitigated after distillation, safety enhancement further reduces the harmfulness of responses.
@@
Safety-enhanced Models 83.12% 66.92% 65.88% 1.55%
Table 3: Overall evaluation metrics of the base model series, distilled model series, and safety-enhanced model series across two tasks. Higher ACC, RR-1 and RR-2 are indicative of better performance, whereas lower HR is preferable. The optimal values under the current metric are highlighted bold.
Following safety enhancement on the distilled models, the accuracy on the risk content identification task increases by 9.12% compared to the distilled models. On the refusal to answer task, the RR-1 and RR-2 metrics improve by 7.56% and 15.23%, respectively, while the HR metric further decreases by 1.58%, achieving comprehensive improvements in safety capabilities. Moreover, the safety-enhanced models not only outperform the distilled models but also surpass the base models, which exhibit slightly stronger safety capabilities: on the risk content identification task, the accuracy increases by 6.95% compared to the base models; on the refusal to answer task, the RR-1 and RR-2 metrics improve by 6.27% and 10.30%, respectively, while the HR metric decreases significantly by 4.03%, demonstrating a complete superiority in safety capabilities.
In summary, the safety-enhanced models demonstrate significantly superior safety capabilities compared to both the non-enhanced distilled models and the non-distilled base models. At the same time, the reasoning capabilities of the safety-enhanced models remain stable, with no significant decline observed, and even improvements in some benchmark tests. These results robustly validate the effectiveness of the safety enhancement strategy.
@@ 5 Conclusion
In response to widespread societal concerns regarding the safety issues of DeepSeek-R1, this paper conducts a systematic safety evaluation of the entire R1 series of distilled models and quantitatively analyzes the negative impact of the distillation process on the safety capabilities of the non-distilled base models. Building on this analysis, the paper further implements targeted safety enhancement measures for all distilled models to rapidly improve their safety performance. Comprehensive evaluation results demonstrate that the safety-enhanced DeepSeek-R1 models achieve a significant safety improvement, even surpassing the base models with stronger safety capabilities, while exhibiting no notable negative impact on reasoning performance.

### v2

@@ 4.3 Reasoning Capability Evaluation
Figure 13: Comparison of DeepSeek-R1 models and safety enhancement models on reasoning benchmarks.
We further evaluate the reasoning capabilities of the safety-enhanced models to investigate whether our safety enhancement methodology compromises their reasoning performance. Figure 13 presents a comparative analysis of reasoning capabilities before and after safety enhancement. As evidenced by the results, safety enhancement does not significantly impair the reasoning abilities of the models; on the contrary, it concurrently enhances safety while improving the reasoning performance of certain models. The distilled models exhibit varied performance on two mathematical reasoning benchmarks: In the MATH-500 test, DeepSeek-R1-Distill-Llama-8B shows a maximum decline of 3.1% in reasoning capability, while DeepSeek-R1-Distill-Qwen-14B achieves a maximum improvement of 1.1%; in the GPQA test, the reasoning capability of DeepSeek-R1 shows a maximum decrease of 1.8%, while DeepSeek-R1-Distill-Qwen-14B demonstrates a maximum increase of 5.0%. Furthermore, in the code benchmark, reasoning capabilities exhibit a marginal reduction, with the decline ranging from 0.3% to 4.1%. Importantly, the observed reductions in code capability across all models are substantially outweighed by the gains in safety performance.
@@ 4.4 Overall Performance
Table 3 presents the overall performance of six DeepSeek-R1 distilled models along with their corresponding base model series and safety-enhanced model series on two Chinese safety evaluation tasks. The aggregate scores for each series were calculated by averaging the scores of their six constituent models. Results demonstrate that in the risk content identification task, the distilled models exhibited a 2.17% accuracy reduction compared to their non-distilled base models. For the refusal-to-answer task, the distilled models showed decreases of 1.29% in RR-1, 4.93% in RR-2, and 2.45% in HR metrics. These findings indicate that while distillation enhances model reasoning capabilities, it concurrently compromises both risk content identification performance and the ability to appropriately refuse answering risky questions while providing responsible responses. Although distillation partially mitigated harmful response rates, our safety enhancement approach achieved further reductions in response harmfulness.
@@
Safety-enhanced Distilled Models 83.12% 66.92% 65.88% 1.55%
Table 3: Overall evaluation metrics of DeepSeek-R1 distilled model series, its corresponding base model series, and safety-enhanced model series across two tasks. Higher ACC, RR-1 and RR-2 are indicative of better performance, whereas lower HR is preferable. The optimal values under the current metric are highlighted bold.
Risk Content Identification Refusal to Answer
ACC RR-1 RR-2 HR
R1 Models 73.63% 60.54% 53.01% 2.78%
Safety-enhanced R1 Models 84.35% 68.93% 67.73% 1.33%
Table 4: Overall evaluation metrics of DeepSeek-R1 model series, and safety-enhanced model series across two tasks. Higher ACC, RR-1 and RR-2 are indicative of better performance, whereas lower HR is preferable. The optimal values under the current metric are highlighted bold.
Table 4 presents a comparative analysis of the overall safety capabilities across all seven DeepSeek-R1 models before and after safety enhancement. In the risk content identification task, the safety-enhanced models demonstrated a 10.72% improvement in accuracy compared to their non-enhanced counterparts. For the refusal-to-answer task, the enhanced models showed significant gains of 18.39% in RR-1 and 14.72% in RR-2 metrics, along with a further 1.45% reduction in HR, achieving comprehensive safety improvement. Furthermore, as evidenced in Table 3, the safety-enhanced models not only outperformed their non-enhanced versions but also surpassed the base models (which originally exhibited slightly stronger safety capabilities) in several aspects: the accuracy in risk content identification improved by 6.95% over base models; the RR-1 and RR-2 metrics increased by 6.27% and 10.30% respectively in refusal-to-answer tasks; and the HR metric decreased by 4.03%, representing a complete outperformance of the base models across all evaluated dimensions.
In summary, the safety-enhanced models demonstrate significantly superior safety capabilities compared to both the non-enhanced models and the non-distilled base models. At the same time, the reasoning capabilities of the safety-enhanced models remain stable, with no significant decline observed, and even improvements in some benchmark tests. These results robustly validate the effectiveness of the safety enhancement strategy.
@@ 5 Conclusion
In response to widespread societal concerns regarding the safety issues of DeepSeek-R1, this paper conducts a systematic safety evaluation of the entire R1 series of distilled models and quantitatively analyzes the negative impact of the distillation process on the safety capabilities of the non-distilled base models. Building on this analysis, the paper further implements targeted safety enhancement measures for all DeepSeek-R1 models to rapidly improve their safety performance. Comprehensive evaluation results demonstrate that the safety-enhanced DeepSeek-R1 models achieve a significant safety improvement, even surpassing the base models with stronger safety capabilities, while exhibiting no notable negative impact on reasoning performance.